### v1

# Testing Hypotheses regarding Covariance and Correlation matrices with the R package CovCorTest

Abstract
In addition to the commonly analyzed measures of location, dispersion measurements such as variance and correlation provide many valuable information. Consequently, they play a crucial role in multivariate statistics, which leads to tests regarding covariance and correlation matrices. Furthermore, also the structure of these matrices leads to important hypotheses of interest, since it contains substantial information about the underlying model. In fact, assumptions regarding the structures of covariance and correlation matrices are often fundamental in statistical modelling and testing. In this context, semi-parametric settings with minimal distributional assumptions and very general hypotheses are essential for enabling manifold usage. The free available package CovCorTest provides suitable tests addressing all aforementioned issues, using bootstrap and similar techniques to achieve good performance, particularly in small samples. Additionally, the package offers flexible specification options for the hypotheses under investigation in two central tests, accommodating users with varying levels of expertise, which results in high flexibility and user-friendliness at the same time. This paper also presents the application of CovCorTest for various issues, illustrated by multiple examples, where the tests are applied to a real-world dataset.
Introduction
Tests for measure of location are widely used in the analysis of multivariate data, such as those arising from repeated measures designs or multiple measurements on the same subject. While expectation vectors are a classical choice for such analyses (Konietschke et al. (2020)), alternative parameters such as quantiles (Baumeister et al. (2024)), relative effects (Rubarth et al. (2022) or Dobler et al. (2020)) and others have been proposed. Withal, measures of dispersion contain substantial information and are, therefore, of central importance in multivariate settings. Especially covariance and correlation matrices are excellent parameters for analyses, which is, among other reasons, because:
  • They are comparably easy to define and to understand.
  • They qualify the relationships between the components, which most other parameters do not take into account.
  • In statistical modeling, covariance and correlation matrices and their structures (e.g. autoregressive or compound symmetry) are essential for understanding and describing the underlying processes.
  • Conditions on the covariance or correlation matrix, such as homogeneity or a special structure/pattern, are among the most frequent assumptions of statistical methods.
  • Finite second moments and therefore, the existence of variances is usually presumed and sufficient to ensure consistency of estimators for the unknown covariance or correlation matrices.
Due to their importance, a variety of statistical procedures have been developed to test hypotheses concerning covariance matrices. Unfortunately, most existing methods rely on restrictive distributional assumptions, such as normality in Box (1953) and Zhong et al. (2017), or specific conditions on the characteristic function in Gupta and Xu (2006), which limits their applicability. Furthermore, several approaches are limited to specific settings, for example, two group settings as in Sakaori (2002) or Omelka and Pauly (2012). Moreover, they typically allow the investigation of only one or two concrete hypotheses, such as equality of covariance matrices (Yang and DeGruttola (2012) or Zhu et al. (2002)) or equality of two correlation matrices (Jennrich (1970)). This also holds for the software packages implementing these procedures, such as the R-packages psych(Revelle (2019)) or CovTestR(Barnard and Young (2018)).
In contrast, the framework introduced in Sattler et al. (2022) allows a broad class of hypotheses, building on methodologies familiar from the analysis of location parameters. At the same time, the less restrictive semi-parametric setting used therein allows broad applicability. Thereto, test statistics are constructed as quadratic forms and Monte-Carlo techniques and resampling approaches are used to obtain the required quantiles and achieve good small sample performance. This approach was further expanded in Sattler and Pauly (2023) to quite general hypotheses regarding correlation matrices, where the focus was the Anova-type-statistic, and additionally, a Taylor-based Monte-Carlo approach was incorporated. The corresponding R package CovCorTest (Sattler et al. (2025)) includes the adequate tests from both papers to provide consistent procedures for investigating covariance and correlation matrices, and moreover also combined tests for equality of covariance and correlation matrices at the same time. Additionally, specific structures of covariance and correlation matrices - such as the Toeplitz form - are of great interest, as they allow for meaningful insights into the underlying model. Besides, specific structures are usual model assumptions, but tests regarding these are often insufficient. Most existing approaches are limited to specific structures such as compound symmetry (Winer et al. (1991) or Votaw (1948)) or first-order autoregressive (McKeown and Johnson (1996)) or focus exclusively on either covariance or correlation matrices (Steiger (1980)). Conversely, the recent work by Sattler and Dobler (2024) introduces a unified framework for testing a large class of structural hypotheses, known as the linear covariance model respectively the linear correlation model, and even further structures. These tests, which are related to the approaches of Sattler et al. (2022) and Sattler and Pauly (2023), were also integrated into the package to enable comprehensive and coherent testing of various aspects of covariance and correlation matrices in one place.
To facilitate access, the package CovCorTest is available for free from the Comprehensive R Archive Network (CRAN) at https://cran.r-project.org/web/packages=CovCorTest.
It contains five main functions (test_covariance, test_correlation, test_covariance_structure,
test_correlation_structure and test_combined), which are as analogous as possible to make it more user-friendly. This uniform design allows researchers to apply the functions in analogous situations targeting different aspects of the data. Moreover, this approach allows for the investigation of all related hypotheses within a single package, while the coherent construction and methodology enhance the comparability of the results. Furthermore, to support users with varying levels of experience, the covariance and correlation tests allow not only the selection of predefined hypotheses, but also the specification of custom hypotheses by providing a hypothesis matrix and the corresponding vector. Since for the remaining tests, either only one hypothesis is possible (test_combined) or only predefined structures can be tested, no such differentiation is necessary there.
Because all underlying approaches and tests are based on empirical covariance and correlation matrices, our functions should only be used for metrical data, although further assumptions, like being a repeated measure design, are not required.
In the subsequent section, we will introduce the statistical multivariate model, the investigated null hypotheses, together with the underlying assumptions. Afterwards, we present the implemented inference procedures, followed by the applications from the R package CovCorTest illustrated by some examples using a real-world dataset. After all, a conclusion and outlook complete the paper.
Statistical Model and Hypotheses
We consider a general multigroup setting with $a\in{\mathbb{N}}$ groups, where each group $i=1,...,a$ contains $n_{i}\in{\mathbb{N}}$ observations. The total number of observations is denoted by $N=\sum_{i=1}^{a}n_{i}$, allowing for potentially unbalanced designs in which group sizes may differ.
Then, a general semi-parametric model is used, defined by independent $d$-dimensional random vectors
$\boldsymbol{X}_{ik}=\boldsymbol{\mu}_{i}+\boldsymbol{\epsilon}_{ik},$
[…]
  • (A1)
    $\frac{n_{i}}{N}\to\kappa_{i}\in(0,1]$, for $\min(n_{1},\dots,n_{a})\to\infty$
  • (A2)
[…]
  • (A3)
    $\operatorname{Cov}(\boldsymbol{\epsilon}_{i1})=\boldsymbol{V}_{i}=(V_{ijk})_{j,k\leq d}\geq 0$
  • (A4)
    $\operatorname{Var}(\boldsymbol{\epsilon}_{i1\ell})=\sigma_{i\ell}>0$,  $\ell=1,...,d$
where (A4) is only required for procedures involving correlation matrices since it guarantees the existence of $\boldsymbol{R}_{i}=\operatorname{Corr}(\boldsymbol{\epsilon}_{i1})$. Additionally, since correlations are only meaningful for at least two components, we assume $d\geq 2$.
Since both covariance and correlation matrices are symmetric, it is natural to consider an appropriate vectorization to transform these matrices into vectors containing the same information as the original matrix. For the covariance matrices, the so-called half-vectorization operation $\operatorname{vech}$ is used, which extracts the upper-triangular (including diagonal) elements of a $d\times d$ matrix, yielding a $p=d(d+1)/2$ dimensional vector. The upper-half-vectorization $\operatorname{vech}^{-}$, on the contrary, only takes the values above the diagonal into a $p_{-}=d(d-1)/2$ dimensional vector, which is useful for correlation matrices whose diagonal entries are always one and hence contain no information. Now, based on hypothesis matrices $\boldsymbol{C}_{\boldsymbol{v}}\in{\mathbb{R}}^{m_{1}\times p}$ respectively $\boldsymbol{C}_{\boldsymbol{r}}\in{\mathbb{R}}^{m_{2}\times p_{-}}$ and vectors $\boldsymbol{\zeta}_{\boldsymbol{v}}\in{\mathbb{R}}^{m_{1}}$, $\boldsymbol{\zeta}_{\boldsymbol{r}}\in{\mathbb{R}}^{m_{2}}$ general linear hypotheses can be formulated. These hypotheses are defined based on the pooled covariance vector $\boldsymbol{v}=(\operatorname{vech}(\boldsymbol{V}_{1})^{\top},\dots,\operatorname{vech}(\boldsymbol{V}_{a})^{\top})^{\top}$ respectively the pooled correlation vector $\boldsymbol{r}=(\operatorname{vech}^{-}(\boldsymbol{R}_{1})^{\top},\dots,\operatorname{vech}^{-}(\boldsymbol{R}_{a})^{\top})^{\top}$ as
$\mathcal{H}_{0}^{\boldsymbol{v}}:\boldsymbol{C}_{\boldsymbol{v}}\boldsymbol{v}=\boldsymbol{\zeta}_{\boldsymbol{v}}\quad\text{ and}\quad\mathcal{H}_{0}^{\boldsymbol{r}}:\boldsymbol{C}_{\boldsymbol{r}}\boldsymbol{r}=\boldsymbol{\zeta}_{\boldsymbol{r}}.$ (1)
Popular hypotheses included in this model, among many others, are
  • Equal covariance matrices: $\boldsymbol{V}_{1}=...=\boldsymbol{V}_{a}$
  • Equal traces of covariance matrices: $\operatorname{tr}(\boldsymbol{V}_{1})=...=\operatorname{tr}(\boldsymbol{V}_{a})$
  • Equal equal variances $(\boldsymbol{V}_{1})_{11}=...=(\boldsymbol{V}_{1})_{dd}$
  • Equal correlation matrices: $\boldsymbol{R}_{1}=...=\boldsymbol{R}_{a}$
  • Equal correlated components: $\operatorname{vech}^{-}(\boldsymbol{R}_{1})_{1}=...=\operatorname{vech}^{-}(\boldsymbol{R}_{1})_{p_{-}}$
  • Uncorrelated components: $\operatorname{vech}^{-}(\boldsymbol{R}_{1})={\bf 0}_{p_{-}}$
  • Trace of the covariance matrix has a special value $\gamma\in{\mathbb{R}}_{+}$: $\operatorname{tr}(\boldsymbol{V}_{1})=\gamma$
  • Specific covariance matrix $\boldsymbol{V}\in{\mathbb{R}}^{d\times d}$: $\boldsymbol{V}_{1}=\boldsymbol{V}$
Note that the same hypothesis can often be expressed in multiple ways through different choices of $\boldsymbol{C}$ and $\boldsymbol{\zeta}$; see Sattler and Zimmermann (2023); Sattler and Rosenbaum (2025) for a detailed discussion. In the case of $\boldsymbol{\zeta}={\bf 0}$, unique projection matrices exist to check the underlying hypotheses, while in general, the considered hypothesis matrices do not have to be square, idempotent or symmetrical. Based for example on the centering matrix $\boldsymbol{P}_{d}=\boldsymbol{I}_{d}-{\bf 1}_{d\times d}/d$ and the Kronecker product, this leads to
$(\boldsymbol{P}_{a}\otimes\boldsymbol{I}_{p})\boldsymbol{v}={\bf 0}_{ap}\quad\Leftrightarrow\quad\boldsymbol{V}_{1}=...=\boldsymbol{V}_{a}\quad\text{resp.}\quad(\boldsymbol{P}_{a}\otimes\boldsymbol{I}_{p_{-}})\boldsymbol{r}={\bf 0}_{ap_{-}}\quad\Leftrightarrow\quad\boldsymbol{R}_{1}=...=\boldsymbol{R}_{a}$
or
$\boldsymbol{P}_{p_{-}}\boldsymbol{r}={\bf 0}_{p_{-}}\quad\Leftrightarrow\quad\operatorname{vech}^{-}(\boldsymbol{R}_{1})_{1}=...=\operatorname{vech}^{-}(\boldsymbol{R}_{1})_{p_{-}}$
and similar for the other hypotheses.
In addition to these general hypotheses, structural hypotheses on the form of the covariance or correlation matrices are of substantial interest. For covariance matrices, common structures include
  • Compound symmetry
  • Diagonality
  • First-order autoregressive (AR(1))
  • Sphericity
  • Toeplitz structure,
see Kincaid (2005) for more details regarding these structures.
For correlation matrices, analogous—but typically fewer—structures exist, which often arise as heterogeneous versions of other structures such as heterogenous autoregressive, heterogeneous compound symmetry or heterogenous Toeplitz.
Except for the autoregressive structure and all its versions, all of these covariance and correlation structures can be expressed within the general hypothesis framework using appropriate choices of $\boldsymbol{C}$ and $\boldsymbol{\zeta}$, as shown in Sattler and Dobler (2024). This highlights the versatility and generality of the presented model in capturing also a wide range of relevant covariance and correlation structures.
Inference
To evaluate the hypotheses formulated in the previous section, we develop inference procedures based on the asymptotic distribution of vectorized covariance and correlation estimates.
For our hypotheses regarding covariance matrices we consider the vectorized empirical group covariances $\widehat{\boldsymbol{v}}_{i}:=\operatorname{vech}\left((n_{i}-1)^{-1}\sum_{i=1}^{n_{i}}(\boldsymbol{X}_{ik}-\overline{\boldsymbol{X}}_{i})(\boldsymbol{X}_{ik}-\overline{\boldsymbol{X}}_{i})^{\top})\right)$ where $\overline{\boldsymbol{X}}_{i}$ is the sample mean of group $i$. Stacking these vectors across all $a$ groups yields the pooled vector $\widehat{\boldsymbol{v}}=(\widehat{\boldsymbol{v}}_{1}^{\top},...,\widehat{\boldsymbol{v}}_{a}^{\top})^{\top}$. Then, for developing a statistical test, we use that under assumption (A2) and (A3) for $i=1,\dots,a$, the group-wise covariance vectors satisfy
$\sqrt{N}(\widehat{\boldsymbol{v}}_{i}-\boldsymbol{v}_{i})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}{\mathcal{N}_{p}\left({\bf 0}_{p},\kappa_{i}^{-1}\cdot\boldsymbol{\Sigma}_{i}\right)}$
with $\boldsymbol{\Sigma}_{i}=\operatorname{Cov}(\operatorname{vech}(\boldsymbol{\epsilon}_{i1}\boldsymbol{\epsilon}_{i1}^{\top}))$. Consequently, under the null hypothesis $\mathcal{H}_{0}^{\boldsymbol{v}}$ and (A1) it holds
$\sqrt{N}(\boldsymbol{C}_{\boldsymbol{v}}\widehat{\boldsymbol{v}}-\boldsymbol{\zeta}_{\boldsymbol{v}})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}{\mathcal{N}_{m_{1}}\left({\bf 0}_{m_{1}},\boldsymbol{C}_{\boldsymbol{v}}\boldsymbol{\Sigma}\boldsymbol{C}_{\boldsymbol{v}}^{\top}\right)},$
with the block-diagonal covariance matrix $\boldsymbol{\Sigma}=\bigoplus_{i=1}^{a}{\kappa_{i}}^{-1}\cdot\boldsymbol{\Sigma}_{i}$.
In the same way, $\widehat{\boldsymbol{r}}_{i}$ denotes the empirical counterpart of the upper-half-vectorized correlation matrix of group $i$ and $\widehat{\boldsymbol{r}}$ is the corresponding pooled vector. Following the results in Sattler and Pauly (2023) (where also the exact form can be found), there exists a continuous matrix-valued function $\boldsymbol{M}\in{\mathbb{R}}^{p_{-}\times p}$ such that
$\sqrt{n_{i}}(\widehat{\boldsymbol{r}}_{i}-\boldsymbol{r}_{i})=\boldsymbol{M}(\boldsymbol{v}_{i},\boldsymbol{r}_{i})\cdot\sqrt{n_{i}}(\widehat{\boldsymbol{v}}_{i}-\boldsymbol{v}_{i})+\scriptstyle\mathcal{O}{}_{P}(1).$ (2)
Using this result and the previous asymptotic expansion, we obtain the joint asymptotic distribution under $\mathcal{H}_{0}^{\boldsymbol{r}}$ as
$\sqrt{N}(\boldsymbol{C}_{\boldsymbol{r}}\widehat{\boldsymbol{r}}-\boldsymbol{\zeta}_{\boldsymbol{r}})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}_{m_{2}}\left({\bf 0}_{m_{2}},\boldsymbol{C}_{\boldsymbol{r}}\bigoplus\nolimits_{i=1}^{a}{\kappa_{i}}^{-1}\boldsymbol{\Upsilon}_{i}\boldsymbol{C}_{\boldsymbol{r}}^{\top}\right)=\mathcal{N}_{m_{2}}({\bf 0}_{m_{2}},\boldsymbol{C}_{\boldsymbol{r}}\boldsymbol{\Upsilon}\boldsymbol{C}_{\boldsymbol{r}}^{\top}),$
where we use $\boldsymbol{\Upsilon}_{i}:=\boldsymbol{M}(\boldsymbol{v}_{i},\boldsymbol{r}_{i})\boldsymbol{\Sigma}_{i}\boldsymbol{M}(\boldsymbol{v}_{i},\boldsymbol{r}_{i})^{\top}$.
Based on these limit distributions, to test our multivariate hypotheses, we use the ATS (Anova-type-statistic), initially developed in Brunner et al. (1997), which is given in this case through:
$ATS_{\boldsymbol{v}}$ $={N}\left[\boldsymbol{C}_{\boldsymbol{v}}(\widehat{\boldsymbol{v}}-\boldsymbol{\zeta}_{\boldsymbol{v}})\right]^{\top}\left[\boldsymbol{C}_{\boldsymbol{v}}(\widehat{\boldsymbol{v}}-\boldsymbol{\zeta}_{\boldsymbol{v}})\right]/\operatorname{tr}(\boldsymbol{C}_{\boldsymbol{v}}\widehat{\boldsymbol{\Sigma}}\boldsymbol{C}_{\boldsymbol{v}}^{\top}),$ (3)
$ATS_{\boldsymbol{r}}$ $={N}\left[\boldsymbol{C}_{\boldsymbol{r}}(\widehat{\boldsymbol{r}}-\boldsymbol{\zeta}_{\boldsymbol{r}})\right]^{\top}\left[\boldsymbol{C}_{\boldsymbol{r}}(\widehat{\boldsymbol{r}}-\boldsymbol{\zeta}_{\boldsymbol{r}})\right]/\operatorname{tr}(\boldsymbol{C}_{\boldsymbol{r}}\widehat{\boldsymbol{\Upsilon}}\boldsymbol{C}_{\boldsymbol{r}}^{\top}).$ (4)
Here for the unknown covariance, we use the estimator $\widehat{\boldsymbol{\Sigma}}:=\bigoplus_{i=1}^{a}{N}/{n_{i}}\cdot\widehat{\boldsymbol{\Sigma}}_{i}$ , where
$\widehat{\boldsymbol{\Sigma}}_{i}=\frac{1}{n_{i}-1}\sum\limits_{k=1}^{n_{i}}\left[\operatorname{vech}\left(\widetilde{\boldsymbol{X}}_{ik}\widetilde{\boldsymbol{X}}_{ik}^{\top}-\sum\limits_{\ell=1}^{n_{i}}\frac{\widetilde{\boldsymbol{X}}_{i\ell}\widetilde{\boldsymbol{X}}_{i\ell}^{\top}}{n_{i}}\right)\right]\left[\operatorname{vech}\left(\widetilde{\boldsymbol{X}}_{ik}\widetilde{\boldsymbol{X}}_{ik}^{\top}-\sum\limits_{\ell=1}^{n_{i}}\frac{\widetilde{\boldsymbol{X}}_{i\ell}\widetilde{\boldsymbol{X}}_{i\ell}^{\top}}{n_{i}}\right)\right]^{\top}.$
is based on $\widetilde{\boldsymbol{X}}_{ik}:=\boldsymbol{X}_{ik}-\overline{\boldsymbol{X}}_{i\cdot}$, the centered version of observation $k$ in group $i$. Based on this, for the correlation we get a plug-in estimator through
$\widehat{\boldsymbol{\Upsilon}}:=\bigoplus_{i=1}^{a}\frac{n_{i}}{N}\widehat{\boldsymbol{\Upsilon}}_{i},\quad\widehat{\boldsymbol{\Upsilon}}_{i}=\boldsymbol{M}(\widehat{\boldsymbol{v}}_{i},\widehat{\boldsymbol{r}}_{i})\widehat{\boldsymbol{\Sigma}}_{i}\boldsymbol{M}(\widehat{\boldsymbol{v}}_{i},\widehat{\boldsymbol{r}}_{i})^{\top}.$
There exists also alternative versions of the ATS, which, for example, allow the approximation through an F-distribution (see e.g. Brunner et al. (2019) or Thiel et al. (2024)). However, for the statistics defined in (3) and (4), the asymptotic null distribution is known, enabling valid inference without stringent distributional assumptions.
Theorem 0.3.1
1
  • a)
    Under Assumption (A1)-(A3) and the null hypothesis $\mathcal{H}_{0}^{v}:\boldsymbol{C}_{\boldsymbol{v}}\boldsymbol{v}=\boldsymbol{\zeta}_{\boldsymbol{v}}$, the $ATS_{\boldsymbol{v}}$ defined by (3) has, asymptotically, a “weighted $\chi^{2}$-distribution”. That is,
    $ATS_{\boldsymbol{v}}\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\sum_{\ell=1}^{m_{1}}\lambda_{\ell}B_{\ell},$
    where $B_{\ell}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\chi_{1}^{2}$ and $\lambda_{\ell},\ell=1,\dots,m_{1},$ are the eigenvalues of $(\boldsymbol{C}_{\boldsymbol{v}}\boldsymbol{\Sigma}\boldsymbol{C}_{\boldsymbol{v}}^{\top})/\operatorname{tr}(\boldsymbol{C}_{\boldsymbol{v}}{\boldsymbol{\Sigma}}\boldsymbol{C}_{\boldsymbol{v}}^{\top})$.
  • b)
    Under Assumption (A1)-(A4) and the null hypothesis $\mathcal{H}_{0}^{r}:\boldsymbol{C}_{\boldsymbol{r}}\boldsymbol{r}=\boldsymbol{\zeta}_{\boldsymbol{r}}$, the $ATS_{\boldsymbol{r}}$ defined by (4) has, asymptotically, a “weighted $\chi^{2}$-distribution”. That is,
    $ATS_{\boldsymbol{r}}\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\sum_{\ell=1}^{m_{2}}\lambda_{\ell}B_{\ell},$
    where $B_{\ell}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\chi_{1}^{2}$ and $\lambda_{\ell},\ell=1,\dots,m_{2},$ are the eigenvalues of $(\boldsymbol{C}_{\boldsymbol{r}}\boldsymbol{\Upsilon}\boldsymbol{C}_{\boldsymbol{r}}^{\top})/\operatorname{tr}(\boldsymbol{C}_{\boldsymbol{r}}{\boldsymbol{\Upsilon}}\boldsymbol{C}_{\boldsymbol{r}}^{\top})$.
A Monte-Carlo technique is one useful approach to obtain the required quantiles of these weighted sums for this statistic. For the covariance matrix, with our consistent covariance estimator $\widehat{\boldsymbol{\Sigma}}$, we get estimators $\hat{\lambda}_{1},...,\hat{\lambda}_{m_{1}}$ for the unknown eigenvalues. With these weights we generate appropriate realisations $B_{1},...,B_{m_{2}}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\chi_{1}^{2}$. These realizations and our estimators allow calculating a weighted sum $\sum\limits_{\ell=1}^{m_{2}}\hat{\lambda}_{\ell}B_{\ell}$, which converges in distribution against the above limit distribution. Repeating this procedure many times allows to calculate the required Monte-Carlo quantiles, and with an analogous procedure for the correlation for $\alpha\in(0,1)$, we get asymptotic level $\alpha$ tests through
$\begin{array}[]{lrl}&\varphi_{\boldsymbol{v}}^{MC}&=11\left(ATS_{\boldsymbol{v}}>q_{\boldsymbol{v},1-\alpha}^{MC}\right),\\ &\varphi_{\boldsymbol{r}}^{MC}&=11\left(ATS_{\boldsymbol{r}}>q_{\boldsymbol{r},1-\alpha}^{MC}\right).\end{array}$
Another technique for obtaining quantiles that requires more computational effort but usually yields better small-sample performance is a bootstrap approach. For hypotheses regarding covariance matrices, the parametric bootstrap mimics the limit distribution and afterwards the structure of the test statistic by generating bootstrap vectors $\boldsymbol{Z}_{i1}^{\star},...,\boldsymbol{Z}_{in_{i}}^{\star}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}_{p}\left({\bf 0}_{p},\widehat{\boldsymbol{\Sigma}}_{i}\right),$ based on given realizations $\boldsymbol{X}_{i1},...,\boldsymbol{X}_{in_{i}}$ with estimators $\widehat{\boldsymbol{\Sigma}}_{i}$. Then $\widehat{\boldsymbol{\Sigma}}_{i}^{\star}$, the empirical covariance matrix of the bootstrap sample $\boldsymbol{Z}_{i1}^{\star},...,\boldsymbol{Z}_{in_{i}}^{\star}$ and $\widehat{\boldsymbol{\Sigma}}^{\star}:=\bigoplus_{i=1}^{a}{N}/{n_{i}}\cdot\widehat{\boldsymbol{\Sigma}}_{i}^{\star}$ are calculated. Following the bootstrap realization $ATS_{\boldsymbol{v}}^{\star}=N[\boldsymbol{C}\overline{\boldsymbol{Z}}^{\star}]^{\top}[\boldsymbol{C}\ \overline{\boldsymbol{Z}}^{\star}]\big{/}\operatorname{tr}(\boldsymbol{C}\widehat{\boldsymbol{\Sigma}}^{\star}\boldsymbol{C}^{\top})$ is calculated, using the bootstrap means $\overline{\boldsymbol{Z}}_{i}^{\star}=n_{i}^{-1}\sum_{k=1}^{n_{i}}\boldsymbol{Z}_{i}k^{\star}$ and $\overline{\boldsymbol{Z}}^{\star}=({\overline{\boldsymbol{Z}}_{1}^{\star}}^{\top},...,{\overline{\boldsymbol{Z}}_{a}^{\star}}^{\top})^{\top}$. Then based on a large number of repetitions, the bootstrap quantile $q_{\boldsymbol{v},1-\alpha}^{\star}$ is determined.
In a similar way, by replacing $\widehat{\boldsymbol{\Sigma}}$ and $\widehat{\boldsymbol{\Sigma}}_{i}^{\star}$ through $\widehat{\boldsymbol{\Upsilon}}$ and $\widehat{\boldsymbol{\Upsilon}}_{i}^{\star}$ the bootstrap quantile for the correlation $q_{\boldsymbol{r},1-\alpha}^{\star}$ can be determined. This allows formulating asymptotic bootstrap tests for the respective hypothesis and $\alpha\in(0,1)$ by
$\begin{array}[]{lrl}&\varphi_{\boldsymbol{v}}^{\star}&=11\left(ATS_{\boldsymbol{v}}>q_{\boldsymbol{v},1-\alpha}^{\star}\right),\\ &\varphi_{\boldsymbol{r}}^{\star}&=11\left(ATS_{\boldsymbol{r}}>q_{\boldsymbol{r},1-\alpha}^{\star}\right).\end{array}$
For the investigation of hypotheses regarding correlations, even one further approach exists to obtain the required quantiles. It is called a Taylor-based Monte-Carlo technique and can be seen as a mixture between Monte-Carlo and bootstrap, using a refined version of the expansion in equation (2) to simulate the asymptotic distribution directly. Further details are provided in Sattler and Pauly (2023).
Remark 0.3.1
Despite its high flexibility, the linear hypotheses introduced in (1) did not allow the formulation of all structures for covariance and correlation matrices. In particular, the autoregressive structure and all of their versions include some proportionality between the matrix components that cannot be captured through linear constraints on the vectorized covariance or correlation matrices. To accommodate such nonlinear structures, we consider transformations of the parameter vectors. Therefore, a kind of differentiable transformation $\boldsymbol{f}$ of the vectorized covariance resp. $\boldsymbol{g}$ of the correlation is required. Therewith, the hypothesis can be formulated through $\mathcal{H}_{0}^{\boldsymbol{v},s}:\widetilde{\boldsymbol{C}}_{\boldsymbol{v}}\boldsymbol{f}(\boldsymbol{v})=\widetilde{\boldsymbol{\zeta}}_{\boldsymbol{v}}$ or $\mathcal{H}_{0}^{\boldsymbol{r},s}:\widetilde{\boldsymbol{C}}_{\boldsymbol{r}}\boldsymbol{g}(\boldsymbol{r})=\widetilde{\boldsymbol{\zeta}}_{\boldsymbol{r}}$. By replacing the vectorized covariance resp. correlation by their transformed pendant, and adding the corresponding Jacobian matrix, the above-mentioned tests can be adapted. A detailed discussion, along with a useful transformation, referred to as the subdiagonal-mean-ratio, is provided in Sattler and Dobler (2024).
Although these tests are very versatile through the less restrictive assumptions and the general class of hypotheses, they do not by themselves yield insight into the specific reason a hypothesis is rejected — e.g., whether the differences arise due to variance differences, dependence structure, or both. Since correlation matrices reflect the dependency structure but contain no information about the components’ variances, but covariance matrices contain no standardized dependency measure, both on their own are not the best solution. To address this issue, Sattler and Pauly (2023) proposes a combined test for simultaneously assessing group differences in both covariance and correlation. This approach is based on the concept of Multiple Contrast Testing (see, e.g., Bretz et al. (2001)) and is applicable when comparing two groups ($a=2$). In case of rejection of equal covariance matrices, it also provides information on whether the dependency, the variance, or both are responsible for it. This test is based on a vector-value test statistic
$\boldsymbol{T}=\begin{pmatrix}T_{1}\\ \vdots\\ T_{p}\end{pmatrix}=\sqrt{N}\left(\begin{pmatrix}\widehat{V}_{111}\\ \widehat{V}_{122}\\ \vdots\\ \widehat{V}_{1dd}\\ \widehat{\boldsymbol{r}}_{1}\end{pmatrix}-\begin{pmatrix}\widehat{V}_{211}\\ \widehat{V}_{222}\\ \vdots\\ \widehat{V}_{2dd}\\ \widehat{\boldsymbol{r}}_{2}\end{pmatrix}\right)$
for which it holds under the $\mathcal{H}_{0}^{\boldsymbol{v}}:\boldsymbol{V}_{1}=\boldsymbol{V}_{2}$ the convergence $\boldsymbol{T}\stackrel{{\scriptstyle\mathcal{D}}}{{\to}}\mathcal{N}_{p}({\bf 0}_{p},\boldsymbol{\Gamma})$, while the explicit form of the asymptotic covariance matrix $\boldsymbol{\Gamma}$ can be found in Sattler and Pauly (2023).
To calculate the required quantiles to use this test statistic, a simulation-based method similar to the Taylor-based Monte Carlo approach is employed to mimic the limit distribution. To this aim for both groups $i=1,2$ independent realisations $\boldsymbol{\Lambda}_{i}\sim\mathcal{N}_{p}({\bf 0}_{p},\widehat{\boldsymbol{\Sigma}}_{i})$ are generated, and subsequently transformed to $\boldsymbol{\Lambda}_{i}^{Tay}$. The extensive formula for the used transformation and their derivation is also part of the online supplement in Sattler and Pauly (2023). From this, we compute $\boldsymbol{T}^{1,Tay}=\sqrt{N}(n_{1}^{-1/2}\boldsymbol{\Lambda}_{1}^{Tay}-n_{2}^{-1/2}\boldsymbol{\Lambda}_{2}^{Tay})$ and repeat this $B\in{\mathbb{N}}$ times, allowing to derive empirical $\beta$ quantiles for each component, denoted as $q_{\ell,\beta}$. Subsequent, the local $\beta$ has to be determined in a way, that it controls the family-wise type-I-error rate by $\alpha$, which is assured through the approach of Munko et al. (2024) by using
${\widetilde{\beta}=\max\left(\beta\in\mathcal{B}\Big{\lvert}\sum\limits_{b=1}^{B}\max\limits_{\ell=1,...,p}\frac{\left(1-11\left(q_{\ell,\frac{\beta}{2}}\leq T_{\ell}^{b,Tay}\leq q_{\ell,1-\frac{\beta}{2}}\right)\right)}{B}\leq\alpha\right)}$
with =$\mathcal{B}=\left\{0,\frac{1}{B},...,\frac{B-1}{B}\right\}$.
An asymptotic level-$\alpha$ test of the null hypothesis of equal covariance matrices is then defined through
$\max\left[\max\limits_{\ell=1,...,d}\left(1-11\left(q_{\ell,\frac{\widetilde{\beta}}{2}}\leq T_{\ell}\leq q_{\ell,1-\frac{\widetilde{\beta}}{2}}\right)\right),\max\limits_{\ell=d+1,...,p}\left(1-11\left(q_{\ell,\frac{\widetilde{\beta}}{2}}\leq T_{\ell}\leq q_{\ell,1-\frac{\widetilde{\beta}}{2}}\right)\right)\right],$
while the arguments resulting in rejection contain further information. If the first argument is equal to one, this means that the variances differ, while a second argument of one implies that the correlation differs. This also leads to three p-values: one for the variances, one for the correlation, and one global p-value, which is the maximum of the other two. In contrast to all other considered tests, this combined test of variance and correlation is only defined for $a=2$ groups.
Software and examples
In this section, we will analyse datasets regarding different hypotheses to illustrate the usage of the testing procedures implemented in the CovCorTest package. Thereto, first, the syntax of our functions is explained, which enables their application to a real-world dataset. Since most of the implemented tests follow a common construction, their application is consistent and intuitive. Therefore, we begin with a general overview of the syntax before highlighting specific differences. Hereby, especially test_combined differs through its more specific hypothesis and setting, resulting in fewer options.
Syntax
  • data: The dataset for which the test should be conducted, while two different input formats are possible. One possibility is a list, where each list element is a matrix containing the values from one group as columns. Since each column is an observation, all matrices must have the same number of rows, while the group sample size and, therefore, the number of columns can differ. The second format are all groups together in one large matrix, where the observations are ordered by the group. In this case, an accompanying vector nv has to be provided, containing the group sample size, while for single-group tests, this value is not necessary. Since the structure tests of covariance or correlation matrices are only applicable for one group, the second format is mandatory in this case.
  • hypothesis:
    • For increased user-friendliness, the functions test_covariance, test_correlation,
      test_covariance_structure and test_correlation_structure allow to choose from a list of predefined hypotheses, which depends on the considered test respectively function:
      Covariance matrix: Here, in the case of one group, the options are "equal" (denoting equal covariances between the components), "given-trace", "given-matrix" and "uncorrelated". In the case of multiple groups, the predefined hypotheses are "equal" (now denoting equal covariance matrices in the different groups), "equal-trace" and "equal-diagonals".
      Correlation matrix: For one group, the hypotheses are "equal-correlated", denoting equal correlation between the components and "uncorrelated", which means that all components have a correlation of zero. In the case of multiple groups, the only hypothesis is "equal-correlated", which here denotes the equality of correlation matrices.
      Covariance Structure: For covariance structures, the selectable hypotheses are autoregressive ("autoregressive" or "ar"), first-order autoregressive ("fo-autoregressive" or "fo-ar"), diagonal ("diagonal" or "diag"), sphericity ("sphericity" or "spher"), compound symmetry ("compoundsymmetry" or "cs") and toeplitz ("toeplitz" or "toep").
      Correlation Structure: The predefined structures of the correlation matrices are heterogenous autoregressive ("hautoregressive" or "har"), heterogenous toeplitz ("htoeplitz" or "htoep"), heterogenous compound symmetry ("hcompoundsymmetry", "hcs") and diagonal ("diagonal" or "diag").
    • For test_covariance and test_correlation, more experienced users could specify custom hypothesis through a hypothesis matrix together with a corresponding vector, as introduced in Sattler et al. (2022) or Sattler and Pauly (2023), especially for less standard hypotheses which increase flexibility. Both refer to the corresponding kind of upper triangular vectorization, and unlike in other procedures, the hypothesis matrix here has neither to be a projection matrix nor square.
    • The function test_combined, which jointly assesses the equality of covariance and correlation matrix, is restricted to one specific hypothesis. Consequently, this argument is not available in that context.
  • method: Specifies the method by which the quantile for the test is calculated. With BT, the quantiles are determined through a parametric bootstrap and with MC through a Monte-Carlo technique, both based on Sattler et al. (2022) and Sattler and Pauly (2023). In the latter, there is even another method, a Taylor-based Monte-Carlo approach, which leads to the option Tay for test_correlation. Again, the combined test works only for a Taylor-based Monte-Carlo approach, which makes any options here unnecessary.
  • repetitions: The number of Monte Carlo or bootstrap repetitions used in the approximation of the quantiles. This number substantially influences both, the test’s performance as well as its computation time. Since a small number of repetitions increases the influence of the randomness on the results, a predefined value of 1000 is a good trade-off and substantially smaller numbers are not recommended.
Output
Here we must differentiate between the combined test, and all other implemented test functions.
For the latter ones, the output consists of different parts, which are mostly general and contains more descriptive information about the conducted test, like the investigated null hypothesis (through the label from the list or the hypothesis matrix $\boldsymbol{C}$ and the corresponding vector $\boldsymbol{\zeta}$), the used method for calculating the critical values, the used number of repetitions and the vector containing the group sample sizes. Moreover, the estimated covariance matrix for the test statistic, as well as the value of the test statistic and the p-value, are returned. The print function only gives a short and compact summary of the conducted test, containing the number of groups, the tested hypothesis, the resampling method together with the used number of repetitions, and especially the value of the test statistic and the p-value.
For the combined test, the hypothesis, the number of groups, and the resampling method are fixed, since only one hypothesis can be tested for two groups. Moreover, based on the used Taylor-based Monte-Carlo approach, no covariance estimator of the test statistic is used and, therefore, also not returned. Moreover, the test statistic here is vector-valued and non-standardised, making the components hardly comparable. Since conclusions based on the value of the test statistic are usually not possible here, these values are not printed to avoid misinterpretations. However, this test results not only in one p-value but also in three of them. A \$pvalue_Variances for the local hypothesis of equal variances, a \$pvalue_Correlations for the local hypothesis of equal correlations, and \$pvalue_Total for the global hypothesis of equal covariance matrices (and therefore both, equal correlations and equal variances). All of them, are returned and printed through the function.
Examples
To demonstrate the usage of the main functions and to illustrate typical outputs, we present several example analyses based on real-world data. For this purpose, we use the dataset EEG, which concludes multivariate measurements from a neurological study conducted at the University Clinic of Salzburg, Department of Neurology. This dataset is part of the R-package manova.rm (Friedrich et al. (2019)), and contains $N=160$ patients from both genders, with three different diagnoses of impairments (subjective cognitive complaints (SCC), mild cognitive impairment (MCI), and Alzheimer’s disease (AD)). Thereby, two neurological parameters (z-score of the brain rate and Hjorth complexity) were measured at three different locations of the head: temporal, frontal, and central. This design results in a multivariate factorial structure with multiple response variables and multiple independent groups and was further described in Bathke et al. (2018). In the following sections, we apply the core testing procedures from CovCorTest to this dataset, showcasing the practical implementation and interpretation of results.
Covariance matrix test
Due to the sample sizes, we will check whether there is homogeneity of covariance matrices for women with these diagnoses, leading to $a=3$ unbalanced groups. The test is done two times, using the Monte-Carlo approach and the parametric bootstrap as a comparison, while for the latter, both kinds of specifying the hypothesis of interest are used.
> data("EEGwide", package = "MANOVA.RM")
> vars <- colnames(EEGwide)[1:6]
> SCCFemale <- t(EEGwide[EEGwide\$sex == "F" & EEGwide\$diagnosis == "SCC",vars])
> MCIFemale <- t(EEGwide[EEGwide\$sex == "F" & EEGwide\$diagnosis == "MCI",vars])
> ADFemale <- t(EEGwide[EEGwide\$sex == "F" & EEGwide\$diagnosis == "AD",vars])
> X <- list(SCCFemale, MCIFemale, ADFemale)
> nv <- unlist(lapply(X, ncol))
> set.seed(123)
> test_covariance(X = X, nv = nv, hypothesis = "equal", method = "MC", repetitions = 1000)
     ¯ Covariance Test
 ¯    3  groups
 Hypothesis: ¯¯equal
 Teststatistic value: ¯2.6042
 p-value: ¯ ¯p = 0.024
 p-value computed using Monte-Carlo-technique with B=1000 repetitions
> set.seed(123)
> test_covariance(X = X, nv = nv, hypothesis = "equal", method = "BT", repetitions = 1000)
       ¯ Covariance Test
 ¯        3  groups
 Hypothesis: ¯¯equal
 Teststatistic value: ¯2.6042
 p-value: ¯ ¯p = 0.023
 p-value computed using Bootstrap with B=1000 repetitions
> P3 <- diag(1,3,3) - 1\3
> Hypothesismatrix <- P3%x%diag(1, 6, 6)
> Hypothesisvector <- matrix(0, 18, 1)
> set.seed(123)
> test_covariance(X = X, nv = nv, C = Hypothesismatrix, Xi = Hypothesisvector,
        method = "BT", repetitions = 1000)
       ¯ Covariance Test
 ¯        3  groups
 Hypothesis: ¯¯equal
 Teststatistic value: ¯2.6042
 p-value: ¯ ¯p = 0.023
 p-value computed using Bootstrap with B=1000 repetitions
As expected, since the value of the test statistic is not affected through the resampling method, both values coincide, while we get slightly different p-values. With p-values of 0.024 resp 0.023, the null hypothesis of equal covariance matrices in all groups would be rejected to the usual level of $5\%$.
Correlation matrix test
After comparing the covariance matrices for women with these diagnoses, the correlation matrices are also investigated for equality. In addition to the parametric bootstrap, the test is also conducted for the Taylor-based Monte-Carlo approach, specifically developed for testing vectorized correlation matrices.
> set.seed(123)
> test_correlation(X = X, nv = nv, hypothesis = "equal-correlated", method = "BT",
                repetitions = 1000)
       ¯ Correlation Test
 ¯    3  groups
 Hypothesis: ¯¯equal-correlated
 Teststatistic value: ¯0.7432
 p-value: ¯ ¯p = 0.602
 p-value computed using Bootstrap with B=1000 repetitions
> set.seed(123)
> test_correlation(X = X, nv = nv, hypothesis = "equal-correlated", method = "Tay",
                repetitions = 1000)
       ¯ Correlation Test
 ¯    3  groups
 Hypothesis: ¯¯equal-correlated
 Teststatistic value: ¯0.7432
 p-value: ¯ ¯p = 0.596
 p-value computed using Taylor-based Monte-Carlo-approach with B=1000 repetitions
Here, the p-values are way higher than for comparison of covariance matrices, such that this hypothesis is not rejected for usual $\alpha$ levels. This is not surprising, since the hypothesis of equal correlation matrices is a subset of the hypothesis of equal covariance matrices.
Combined test
Instead of testing the equality of covariances matrices and the equality of correlation matrices for two groups separately, the combined test from Sattler and Pauly (2023) allows us to do both at the same time. This allows for the null hypothesis of equal covariance matrices, to conduct whether the difference lies in the correlations and, therefore, the dependency, the variances, or even both.
> data("EEGwide", package = "MANOVA.RM")
> vars <- colnames(EEGwide)[1:6]
> X <- list(t(EEGwide[EEGwide\$sex == "M" & EEGwide\$diagnosis == "AD",vars]),
                t(EEGwide[EEGwide\$sex == "M" & EEGwide\$diagnosis == "MCI",vars]))
> nv <- unlist(lapply(X, ncol))
> set.seed(123)
> test_combined(X, nv)
       ¯ Combined Test
 p-value-Variances: ¯ ¯     p = 0.418
 p-value-Correlations: ¯ ¯  p = 0.016
 p-value-Total: ¯ ¯         p = 0.016
 p-values computed using Taylor-based Monte-Carlo-approach with B=1000 repetitions
From the resulting three p-values, we see that both covariance matrices differ, while only a difference in the correlation can be conducted, and it was not possible to prove a difference in the variances.
Structure tests
In this analysis, we want to investigate whether the position of the measuring points influences the measured values. If we reject the hypothesis of the covariance matrix being a compound symmetry matrix, it is verified that there is an effect of the position. Moreover, to check whether also the correlations between the measure points are equal, we additionally check for a heterogeneous compound symmetry structure by using the correlation matrix.
> X <- as.matrix(EEGwide[EEGwide\$sex == "W" & EEGwide\$diagnosis == "AD",
     c("brainrate_temporal", "brainrate_frontal","brainrate_central",
     "complexity_temporal","complexity_frontal", "complexity_central")])
> set.seed(123)
> test_covariance_structure(X = X, structure = "compoundsymmetry", method = "MC")
       ¯ Covariance Test
 ¯    one group
 Hypothesis: ¯¯compoundsymmetry
 Teststatistic value: ¯3.055
 p-value: ¯ ¯p = 0.026
 p-value computed using Monte-Carlo-technique with B=1000 repetitions
> set.seed(123)
> test_correlation_structure(X = X, structure = "hcompoundsymmetry", method = "MC")
 ¯ Correlation Test
 ¯    one group
 Hypothesis: ¯¯hcompoundsymmetry
 Teststatistic value: ¯5.5229
 p-value: ¯ ¯p < 1e-3
 p-value computed using Monte-Carlo-technique with B=1000 repetitions
Since both structures are rejected to a usual significance level of $5\%$, not only an influence of the measure points’ location was verified, but also the correlation between them differs.
Conclusion and Outlook
In the presented paper, we illustrate the functionality and scope of the R package CovCorTest, which provides a comprehensive framework for hypothesis testing concerning covariance and correlation matrices in multivariate data analysis. It contains tests for both parameters and different hypotheses in a uniform way, which excels through its applicability to large class of possible hypotheses and less distribution assumptions. Apart from multi-group settings, single groups are also allowed as well as factorial designs.
In the main, it contains functions for five different tests: regarding covariance matrices, correlation matrices, a combined test for equality of covariance and correlation matrices, and tests regarding their corresponding structures. To this aim, quantiles based, among other things, on resampling techniques and Monte-Carlo approaches are used. Although the Monte-Carlo approach leads to the least time-consuming procedure, a parametric bootstrap for covariances and a Taylor-based Monte-Carlo approach for correlations are preferable for smaller sample sizes.
To include actual developments and procedures in this area, we aim to update this package on a regular basis. Planned future developments include further hypotheses (especially additional structures and patterns) and new test statistics, such as multiple contrast tests for hypotheses regarding covariance and correlation matrices. Also, the combined correlation and covariance test could be expanded to accommodate more than two groups, and tests for multiple structures at once, as mentioned in Sattler and Dobler (2024), could be incorporated. Further, the output of a confidence ellipsoid is planned as a further extension, together with the corresponding plot. Also, a graphical user interface (GUI) is possible to simplify the usage, especially for more applied users. Finally, the results from Sattler and Rosenbaum (2025) could be realized in the future, to reduce calculation effort. This, would allow a higher number of bootstrap and Monte-Carlo repetitions and thereby increase the performance for large datasets.
Acknowledgement
Svenja Jedhoff gratefully acknowledges the support from TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (520388526), funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).